Find the present value for the following string of cash flows. The risk free rate is currently 1%. A stock has a beta of 1.28 and the market risk premium is 8%.

Year 1 cash flow: $1,500

Year 2 cash flow: $5,500

Year 3 cash flow: $2,010

cash flows will grow 2% indefinitely after year 3

A.$17,795.56

B.$18,668.82

C.$23,372.44

D.$25,468.65
Answer #1

Required rate=risk free rate+Beta*market risk premium

=1+(1.28*8)=11.24%

Value after year 3=(Year 3 cash flow*Growth Rate)/(Required Return-Growth Rate)

=(2010*1.02)/(0.1124-0.02)

=$22188.31169

Hence present value=Future values and cash flows*Present value of discounting factor(rate%,time period)

=1500/1.1124+5500/1.1124^2+2010/1.1124^3+$22188.31169/1.1124^3

=$23372.44(Approx).